### v1

# Attosecond-resolved probing of recolliding electron wave packets in liquids and aqueous solutions

Abstract
High-harmonic spectroscopy (HHS) in liquids promises real-time access to ultrafast electronic dynamics in the native environment of chemical and biological processes. While electron recollision has been established as the dominant mechanism of high-harmonic generation (HHG) in liquids, resolving the underlying electron dynamics has remained elusive. Here we demonstrate attosecond-resolved measurements of recolliding electron wave packets, extending HHS from neat liquids to aqueous solutions. Using phase-controlled two-colour fields, we observe a linear scaling of the two-colour delay that maximizes even-harmonic emission with photon energy, yielding slopes of 20855 as/eV in ethanol and 12442 as/eV in water, the latter matching ab initio simulations (12548 as/eV). In aqueous salt solutions, we uncover interference minima whose appearance depends on solute type and concentration, arising from destructive interference between solute and solvent emission. By measuring the relative phase of solvent and solute HHG, we retrieve a variation of electron transit time by 11332 as/eV, consistent with our neat-liquid results. These findings establish HHS as a powerful attosecond-resolved probe of electron dynamics in disordered media, opening transformative opportunities for studying ultrafast processes such as energy transfer, charge migration, and proton dynamics in liquids and solutions.
Introduction
Electron dynamics in liquids underpin fundamental processes ranging from radiation-induced damage to photocatalysis and prebiotic molecule formation [1, 2, 3, 4, 5, 6, 7, 8]. While these phenomena unfold on femtosecond timescales, accessing the earliest, attosecond-scale electron motion is essential to understanding how chemical reactivity, charge transfer, and energy redistribution are initiated in complex environments [9, 10, 11, 12, 13, 14]. Over the past two decades, attosecond spectroscopy has revolutionized our understanding of electron dynamics in gases and solids, providing insights into tunneling delays, orbital reshaping, and band-structure evolution in strong laser fields [15, 16, 17, 18, 19].
These breakthroughs have been enabled primarily by HHS – a technique that encodes attosecond-resolved temporal information into the phase and amplitude of the emitted radiation. In gases, this arises from the recollision of a liberated electron with its parent ion, providing insights into tunneling, ionization, and recombination dynamics [20, 16, 21]. In solids, by contrast, HHG stems from interband recombination and intraband current modulation, allowing access to band dynamics, phase transitions, and carrier scattering processes in strong fields [15, 18, 19, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33].Unambiguous determination of the harmonic phase and its mapping to the different components of the electron motion, namely ionization, propagation and recombination, is central towards reconstructing ultrafast electron dynamics in quantum systems.
Yet, despite its promise, obtaining time-resolved information from liquid-phase HHS has remained elusive, with the link between the emitted photon energy and the characteristic time scales of the electron dynamics in the liquid phase unexplored to date. This limitation arises from the unique challenges posed by liquids: high density, strong optical absorption, electron scattering and lack of long-range order, all of which make ultrafast measurements and computations demanding. While ultrafast techniques such as photoelectron spectroscopy and transient X-ray absorption spectroscopy have been extended to bulk liquids to investigate electronic and molecular dynamics [34, 14], attosecond time-resolved measurements have so far remained restricted to a measurement of the photoionization delays of liquid- vs. gas-phase water [9, 12] and a single pump-probe delay of X-ray transient absorption [13].
Recent experiments have successfully extended HHG to the liquid phase [35, 36, 37, 38, 39, 40, 41], identifying electron recollision as a viable mechanism in bulk liquids [39, 42, 43, 44]. Unlike in gases, where electron quantum paths are dominated by laser parameters, the maximum quantum path length in liquids at ambient pressures is limited primarily by the intrinsic mean free path (MFP) of low-energy electrons and is largely independent of laser wavelength, intensity, or pulse duration [39, 42].
However, attoseond-resolved HHS in liquid phase has not been achieved to date, and it has remained unclear how the timing of the recolliding electron wave packets could be measured.
In this work, we introduce a new paradigm for attosecond-resolved spectroscopy of liquids and aqueous solutions. We demonstrate all-optical attosecond interferometry in the liquid phase using two complementary, yet independent, approaches:
- 1.
2.
Passive interferometry, exploiting spectral interference between HHG from solute and solvent species in salt solutions to extract the electron transit times (as illustrated in Fig. 1c).
These methods offer two complementary views on the attosecond dynamics of electron dynamics in liquids with 30 as/eV precision.
Remarkably, these two techniques provide independent and matching verifications of the characteristic timescales of ultrafast electron dynamics in solutions, each with its own advantages.
Rather than being coincidental, we argue that this agreement reflects a weak dependence of the ionization times on the emitted photon energy, in agreement with the quantum-orbit description of HHG.
By establishing a mapping of emitted photon energy to the attosecond dynamics of the electronic wavepacket in chemically relevant environments under ambient conditions, this work expands the domain of attosecond HHS into uncharted, but highly relevant regimes.
The ability to track charge motion with sub-cycle precision in liquids lays the foundation for future studies of charge migration in solution, ultrafast solvation dynamics, and light-induced control of liquid-phase chemical reactions, with potentially far-reaching implications for photochemistry, biophysics, and energy science.
Results
We first discuss the two interferometric techniques, starting with Fig. 1a that illustrates active interferometry. This method relies on HHG driven by a phase-controlled two-colour laser field, composed of a fundamental frequency and a weak second-harmonic (2) component. In contrast to a single-colour laser field, where electron quantum paths in consecutive half cycles of the electric field are symmetric and only odd harmonics are emitted [45], the addition of a weak 2 component breaks this symmetry and leads to the generation of even harmonics with an intensity that oscillates periodically with the phase delay between the two colours.
Owing to the different phases accumulated by the electron wave packet in two consecutive half cycles, each even harmonic maximizes when the amount of symmetry breaking along its quantum path is maximal – encoding information about the emission time of the harmonics into the two-colour delay at which the even harmonic maximizes [20, 18].
As shown in Fig. 1b, we observe clear modulations of all even harmonics, whereby the maxima shift to larger delays as a function of the emitted photon energy . This indicates that higher harmonic orders are emitted later than lower ones, rather than being synchronized. In other words, it reveals that high-harmonic emission from liquids is positively chirped.
In contrast to active interferometry, passive interferometry measures the phase difference between harmonic emissions from two species with different ’s. Figure 1c illustrates this concept. In a solution, the different ’s of solute and solvent lead to different electron transit times in the laser field. As a result, the emitted harmonics from each species acquire distinct phases. When the resulting phase difference, , is close to an odd multiple of , destructive interference occurs at that harmonic order. In atoms and molecules, the phase shift between harmonics emitted by species with different ’s follows the relation , where is the binding-energy difference, is the transit time of electron wavepackets contributing to the photon energy [46] and is the dipole phase difference of the two emitting species [47, 48, 49, 50]. Here, we demonstrate that this principle can be extended to bulk liquids and aqueous solutions within the first plateau (whereas a different result is expected in recently observed second-plateau harmonics that arise from a distinct physical mechanism [51]). Figure 1d shows that the destructive interference condition is fulfilled for a 0.8-M solution of NaCl in water, driven by an 1800 nm field of a peak intensity 2.61013 W/cm2. This enables us to probe how the harmonic phase varies with photon energy and extract attosecond time-resolved information about the underlying electron dynamics.
With the conceptual framework in place, we next examine what each interferometric approach reveals about attosecond electron dynamics in liquids.
Active interferometry: Two-colour field modulation and recollision timing in liquids
The schematic of the experimental setup is shown in Fig. 2a. A phase-controlled two-colour (-) laser field with parallel polarizations of the two colours is generated by frequency doubling an 1800-nm fundamental pulse in a -barium borate (BBO) crystal.
The phase delay between the and 2 components is controlled using a pair of fused-silica wedges. The high harmonics generated by this two-colour field in a liquid flat jet or an interchangeable gas jet are detected in a flat-field imaging spectrometer.
Figure 1b shows a typical harmonic spectrum of liquid ethanol as a function of the two-colour delay. Despite the low relative intensity of the 2 component (0.07 %), the symmetry of the HHG process is significantly broken, resulting in the appearance of even harmonics. These harmonics exhibit periodic intensity modulations as a function of the -2 delay, with a modulation period corresponding to the optical cycle of the 2 field. The black-dashed line connects the delay values that maximize the yield of each even harmonic, indicating a linear dependence of the optimal delay on the harmonic order.
We then extract the two-colour delay that maximizes a given harmonic order by Fourier analysis. As shown in Figure 2b, these delays increase approximately linearly with harmonic order for both liquid water (blue circles) and ethanol (orange circles). A linear fit to these data returns slopes of 208 55 as/eV for liquid ethanol and 124 42 as/eV for liquid water.
The experimental results are further supported by ab-initio time-dependent density-functional-theory (TDDFT) calculations on large water clusters (see methods section for details), which yield a slope of 125 48 as/eV (Fig. 2d).
The magnitude of these slopes is much larger than in the gas-phase measurements, performed under comparable conditions. Due to its higher and longer electron quantum paths, the gas-phase experiment required higher driving intensity ( W/cm2) and a stronger 2 component () to observe a clear modulation of the even harmonics. As shown in Fig. 2c (yellow circles), the extracted slope ( as/eV) is much smaller than in the liquid-phase measurements. This result agrees with calculations based on the SFA for a model system with 10.48 eV, matching that of gas-phase ethanol (orange circles in Fig. 2c). Further experimental and theoretical details are provided in the Methods section.
The most striking result of these measurements is the large linear slope, which reflects a large positive attochirp of harmonic emission from liquids.
To extract additional information about the attosecond electron dynamics, we turn to passive interferometry. By independently extracting the relative phases of high harmonics emitted by different species, we assess whether the large variation of the harmonic emission time observed in active interferometry is a general feature of the liquid-phase HHG mechanism, and whether it persists across different measurement techniques.
Passive interferometry: Single-colour measurements in salt solutions.
b TDDFT calculation of the high-harmonic spectrum of a 1-M NaCl solution in water using an 1800-nm driving pulse with a peak intensity of 21013 W/cm2. c Measured yield of H17 (11.7 eV) as a function of NaCl concentration (circles). The orange dashed line is a fit based on a two-species (H2O and Cl-) interference model. d Variation of the relative phase of emission from Cl- and H2O, extracted from a fit to the interference model (left axis) and variation of the electron transit time (right axis), as a function of the emitted photon energy.
In its simplest implementation, this technique requires two species contributing to the harmonic emission. We realize the two-emitter system with aqueous solutions of NaCl, containing H2O, hydrated cations (Na) and anions (Cl), whereby we drop the subscripts “(aq)” for brevity from hereon. Owing to the significantly higher of Na+ (35.4 eV) compared to liquid H2O (11.67 eV) and Cl- (9.55 eV) [52, 53, 54], the cations do not contribute measurably to high-harmonic generation. This statement is confirmed by the experimental observation that the emitted HHG spectra and especially the position of the interference minimum strongly depend on the nature of the anion (Extended Data Fig., EDF 1), but not on that of the cation (EDF 2).
The concentration of Na+ in H2O serves as the control parameter. Note that we use the relative ’s, referenced to the vacuum level, rather than the energy difference between valence bands and the conduction band of the solution, because (i) the ’s are more accurately known than the band gaps and (ii) the relative ’s also reflect the energy difference of the valence bands because of the common conduction band for solute and solvent (see EDF 5d and 6d), in analogy with negatively doped semiconductors [55].
Figure 3a shows high-harmonic spectra as a function of the NaCl concentration. A pronounced minimum in the yield of H17 (11.7 eV) appears around 1 M, with the inset for 0.8 M illustrating the maximal destructive interference. The relative yields of all harmonics vary systematically with concentration, consistent with a phase-sensitive interference between the harmonic emissions from H2O and Cl-.
To further demonstrate that this interference reflects a genuine microscopic effect, we performed elaborate ab-initio TDDFT simulations combined with molecular dynamics (MD) of the salt solution (see the Methods section for details), with the results shown in Fig. 3b. The calculations reproduce the observed minimum (upshifted to the next higher order), confirming that it arises from the microscopic electronic response as opposed to macroscopic phase-matching and propagation effects, which are not included in our TDDFT simulations.
While the ab-initio simulations confirm the microscopic origin of the harmonic suppression, performing a fine-grained concentration scan is computationally expensive and time-consuming, and can become intractable for either very low, or very high, solute concentrations (requiring either very large simulation cells, or high level theory for properly capturing the interaction between the solute ions, respectively). For the interpretation, we therefore introduce an analytical two-emitter interference model (sketched in Fig. 3a), which offers an intuitive and transparent way of determining the relative phase shifts and emission-amplitude ratios.
To extract the phase shifts between H2O and Cl- from experimental data, we fit each harmonic yield as a function of salt concentration using the two-emitter model. For a given solution containing moles of H2O and moles of Cl- ions, the total intensity at a given photon energy is treated as the coherent sum of the contributions from each species:
where and are the single-species emission intensities from 1 mol of particles, and is their relative phase. Dividing this intensity by the known harmonic yield of an identical pure water sample, gives the normalized yield (Eq. Passive interferometry: Single-colour measurements in salt solutions.), which displays the expected quadratic dependence on the molar ratio (). The parameters and are then retrieved from the measured intensities of HHG from solutions of different concentrations.
Figure 3c shows the fit of the two-emitter model (red dashed line) to the normalized yield of harmonic 17 (H17, 11.7 eV). At low concentrations (up to 1.5 M NaCl) the two-emitter model reproduces the data well (Fig. 3c), while at higher concentrations, the measured harmonic yields lie below the model prediction.
This deviation likely has several underlying causes. First, as the salt concentration increases, the probability for the electron wave packet to scatter off Na+ or Cl- ions rises, since both ions have larger electron scattering cross sections than water. This increased scattering reduces the electron’s mean-free path, suppressing the harmonic yield [39]. Similar effects have been reported for higher-energy electrons ( 65 eV), whose mean-free paths decrease with increasing NaI concentrations above 1 M [56]. Second, the addition of salt also increasingly modifies the electronic structure of the solution (EDF 3), especially at higher concentrations where it also changes the geometric structure (EDF 4) which can be expected to alter the HHG yield.
At low concentrations ( M), the validity of the two-emitter model is further supported by our electronic-structure calculations combined with the MD simulations, shown in EDF 5 and 6. They indicate that the HOMO in salt solutions remains highly localized on individual water molecules and anions (EDF 5). Meanwhile, the LUMO is delocalized over the simulation box. The localization of the HOMO ensures that the harmonic emission can be treated as originating from discrete species with well-defined ’s, consistent with the observed interference behavior and supporting the validity of the method chosen for extracting the harmonic phase shift. The delocalization of the LUMO and higher-lying states validates the treatment of the electron dynamics as taking place in a quasi-continuum.
The two-emitter model is also supported by experiments in which we systematically varied the anion. Replacing Cl- with Br- and I- featuring ’s of 9.55 eV, 8.8 eV and 7.9 eV, respectively, and thus an increasing relative to water, we observe that the suppressed harmonic order moves to lower photon energies (EDF 1) – consistent with the fact that lower harmonic orders correspond to the shorter transit times that are necessary to fulfill the destructive-interference condition. Varying the cation has no significant effect, as expected (EDF 2).
Figure 3d shows the extracted phase as a function of the harmonic order. Assuming that the dipole phase difference depends weakly on the photon energy, the variation of can be converted to the variation of the transit time, shown on the right axis of the figure. The assumption of the weak dependence of on the photon energy, which is expected in the absence of (anti-)resonances as confirmed in Ref. [50], is validated in EDF 7.
b Comparison of the extracted variations of ionization (yellow circles) and emission times (blue circles) with a quantum-orbit calculation (lines) described in the Methods Section, at a peak intensity of 3.31013 W/cm2, for 8.83 eV (bandgap of liquid water). The experimental data has been shifted by constants offsets of 0.7 fs (for ionization time, yellow circles) and 1.9 fs (for the emission time, blue circles), to overlay on the curves of the quantum-orbit calculations. These constant shifts were determined from the mean of the difference of the experimental and quantum-orbit results. The driving wavelength used in all experiments and calculations is 1800 nm.
A comparison of the obtained variation of transit times (Fig. 3d) with the results of the two-colour measurements on liquid water (Fig. 2b), shown in Fig. 4a, demonstrates that the determined time intervals agree within the respective experimental errors. This agreement is remarkable, especially for a dynamical system as complex as liquid water, and reflects the attosecond dynamics of electron recollision.
Discussion and conclusions
The attosecond electron dynamics underlying HHG in liquids can be described by three characteristic times: the ionization time (), the transit time (), and the harmonic emission time (). Each of them is encoded in the phases measured in the present work. As explained above, the relative phase of the harmonic emission from H2O and Cl- gives access to the variation of the transit time and the two-colour delays give access to the variation of [20] with photon energy.
Comparing the variations of and in Fig. 4a, we find that they agree within the experimental error. This, in turn, indicates a negligible dependence of the ionization times on the harmonic order. This conclusion is remarkable in that it agrees with the results from gas-phase experiments and the quantum-orbit analysis of gas-phase HHG [16], which both demonstrate that the variation of the ionization time with harmonic order is negligible compared to the variation of the emission time with harmonic order.
This experimental observation suggests to attempt a comparison of the measured variations of and with a quantum-orbit calculation that uses the band gap of liquid water as the .
Figure 4b shows the results of such a calculation (full lines), superimposed with the experimental data – shifted horizontally since only their variation with photon energy is accessible in the experiment. Strikingly, the variations of both and agree with the quantum-orbit calculations within the experimental errors.
Our results have a number of implications. First, the agreement between active and passive interferometry validates both techniques. It reveals that high-harmonic emission from liquids displays a positive linear chirp, which encodes the attosecond dynamics of the underlying electron quantum dynamics. The common energy dependence of both emission and transit times implies a negligible variation of the ionization times with harmonic order, consistent with the quantum-orbit description of HHG.
Second, they establish a quantitative link between the emitted photon energies and the attosecond-scale variation of the recombination times of the electron wave packets with photon energy in the liquid phase. This link turns liquid-phase HHS into a technique with attosecond temporal resolution, capable of resolving electron dynamics on their natural timescale.
Third, our measurements demonstrate that high-harmonic emission from solvent and solute molecules is fully coherent at low concentrations (up to 1-M NaCl or NaBr) within the sensitivity of our measurements and can be used to retrieve the variation of electron transit times on attosecond scales.
Taken together, these results demonstrate all-optical attosecond spectroscopy of electronic processes in liquids and solutions, which can be developed further to study electron tunneling, charge- and energy-transfer processes and electronic dynamics underlying radiation damage, nucleation, proton transfer or phase transitions, as well as electronic processes induced by strong electromagnetic fields.
Methods
.1 Experimental Methods
The experiments were carried out at the fundamental laser wavelength of 1800 nm, generated using an optical parametric amplifier (HE-TOPAS) pumped by a 800-nm, 1-kHz titanium-sapphire beam with a pulse energy of 6 mJ. For the two-colour experiments, the fundamental beam centred at 1800 nm () of pulse energy 400 µJ and pulse duration 55 fs, was used to generate a second harmonic centred at 900 nm (2) using a 75 µm -barium borate (BBO) crystal. The pulse energy ratio between the and the 2 was varied from 0.03% to 0.7% by tilting the BBO with respect to the polarization axis of the 1800-nm beam. A dual wave plate was used to align the and 2 polarizations, and their relative delay was adjusted via a pair of fused-silica wedges. The beams were then focused onto a liquid flat-sheet target, and the emitted high harmonics were detected with a flat-field imaging spectrometer. The flat jet is created by colliding two cylindrical jets created from two cylindrical nozzles of inner diameters 54 m, resulting in a thickness of 1 m [36, 37, 57].
A gas target was mounted on the same manipulator as the liquid jet, to facilitate gas-phase measurement by mere translation of the stage.
To ensure that the experiment operates in the perturbative regime of the 900-nm component, we verified two conditions: (1) the absence of higher-order components (e.g., 4) in the modulation of the even-harmonic yield, and (2) the out-of-phase oscillations of the even- and odd-harmonics intensities.
The 2-phase component of the even-harmonic modulation was extracted using Fourier-transform analysis. For each harmonic, the intensity modulation as a function of the –2 delay was Fourier transformed, and the phase of the 2 peak was extracted from the Fourier spectrum.
For the concentration scan up to 4-5 M salt concentrations, larger nozzle diameters of 100 m were used, to prevent clogging of the jet. However, as the HHG is generated only from the last few 100 nm [36, 39], a thicker flat-jet 2-3 m, does not significantly modify the spectral features of the emitted HHG. Data recorded with both 100 m nozzles (Fig. 3a) and 54 m nozzles (Fig. 4b), show the same selective suppression of the 11.7-eV, H17 at 0.8 M NaCl.
.2 Theoretical Methods
TDDFT cluster calculations. We performed ab-initio TDDFT simulations for the case of liquids driven by intense laser pulses. These simulations followed the approach extensively developed and described in Ref. [58]. We employed a 54-water-molecule cluster [59] that was driven by intense laser pulses as detailed in the main text. The induced dipole response was extracted by solving the time-dependent Kohn-Sham (KS) TDDFT equations of motion employing the Perdew-Burke-Ernzerhof (PBE) [60] exchange-correlation (XC) functional for the ground state, while neglecting electron-electron interactions (employing the independent particle approximation (IPA)) in the time-dependent simulations, and artificially removing the response of surface-localized states (see details in Ref. [58]). We assumed clamped nuclei, which should be applicable up to tens of femtosecond timescales [61]. Simulations employed a spherically-shaped real-space grid with cartesian coordinates of spacing 0.4 Bohr utilizing the Octopus code [62], and a time-step of 1.9 as for the temporal propagation. Further, we employed a complex absorbing potential to remove reflections from the boundaries similar to the approach in Ref. [58]. From the time-dependent KS states we extracted the dipolar response, (with the time-dependent electron density) and the dipole acceleration, . The HHG spectrum was calculated by Fourier transforming the dipole acceleration with an added super-gaussian temporal window function, : . Simulations were performed for driving fields with varying values of the relative two-colour phase in steps of 0.05. In each case the orientation of the cluster was further angularly-averaged to ensure a limit of an isotropic liquid medium (with 12 orientations averaged per pulse, employing additional symmetries of the laser pulse to perform Euler trapezoidal angular integration, as detailed in refs. [58, 63]). We directly averaged the dipole acceleration in each case such that the actual expression from which the HHG spectra was calculated was of the form , with labeling the angular orientation and the integral performed over the Euler angles. Note that the optical gap obtained in the cluster case of liquid water at the PBE level ( eV) is smaller than the experimental value (at eV), but larger than the value obtained in the supercell simulations described below (at eV, since we artificially removed surface states which slightly increase the gap). From the calculated HHG responses, we extracted the two-colour phase dependence of the harmonic integrated yields, as discussed in the main text, and with an approach identical to the experimental scheme (i.e., directly from the numerical data without further assumptions). This was achieved by Fourier transforming the two-colour phase-dependent harmonic integrated yields to obtain the phase-of-the-phase, which was unwrapped over the even harmonics.
MD simulations. To obtain a selection of atomic positions (snapshots) for the TDDFT supercell calculations described below, we performed thermal sampling of the machine learning potential trained by O’Neill and co-workers [64, 65] on DFT forces and energies computed with the dispersion-corrected revPBE-D3 functional [66, 67]. The dynamics were propagated in i-PI [68], using a stochastic velocity rescaling thermostat [69] to equilibrate to a temperature of 300 K. In order to capture nuclear quantum effects in the thermal distribution, we continued sampling using the path-integral molecular dynamics (PIMD) [70] of 12-bead ring polymers coupled to a PIGLET thermostat [71]. The HHG calculations are highly sensitive to the intermolecular structure, and revPBE-D3 slightly overbinds the liquid [72]. To compensate for this effect, we considered PIMD simulations at a selection of densities, namely 0.90, 0.95, 1.00, and 1.10 g/cm3. Overall, HHG spectra derived from snapshots of pure water at 0.90 g/cm3 corresponded best to the experimental signal.
[…]
1.000 g/cm3, and 1.032 g/cm3, respectively, corresponding to experimental densities at 300 K [73], scaled by a factor of 0.9. For every salt concentration, 50 snapshots were recorded at 5 ps intervals and subsequently fed into the HHG calculations.
Ground-state DFT calculations.
We analyzed the electronic density of states (EDOS) and the band structures of the salt solutions, performing ground-state electronic structure calculations on the simulation snapshots using density-functional theory with the PBE [60] and the PBE0 [74] exchange-correlation functionals. The calculations were performed using the FHI-aims software package [75, 76, 77],
with the default tight settings for the basis sets and real-space quadrature grids. The EDOSs were computed at the -point. The total and species-projected EDOSs, obtained with PBE and PBE0 and averaged over all snapshots, are shown in EDF 3. The HOMO and LUMO orbitals, as well as the band structures for selected individual snapshots are shown in EDF 5 (EDF 6) using the PBE0 (PBE) functionals.
TDDFT supercell simulations. Complementary to the cluster simulations of liquid HHG, we employed a secondary ab-initio approach using a supercell of liquid water with 100 molecules in a cubic unit cell (depending on the salt concentration the number of molecules slightly differs). The approach is similar in spirit to simulations in refs. [78, 61], but employing the MD snapshots derived with the method described above, and with some minor differences. We assumed clamped nuclei during the interaction with the laser pulse, which should be valid on fs timescales (though longer timescale picosecond dynamics and structural rearrangement were considered by averaging over multiple liquid snapshots obtained from the MD simulations as is detailed below). The main logic of employing this approach, that is complementary to the cluster simulations, is that the cluster approach cannot faithfully describe HHG from salt solutions, because the cluster geometry is adopted from previous work, and the clusters are too small to describe a bulk solution. Moreover, multiple MD snapshots are required to obtain a proper electronic structure in such systems [79]. In that respect, the supercell scheme is expected to be more accurate than the less costly cluster simulations (however, we note that due to the high numerical cost of this approach requiring averaging over many MD snapshots and employing large supercells it cannot be employed for simulating the experiments).
With this in mind, simulations in the main text that involve HHG for different salt concentrations in water were performed by solving the TDDFT KS equations of motion fully (for all states in the system) in a periodic cubic supercell, using the Octopus code [62], and employing a real-space grid with -point only k-space integration and 0.4 Bohr spatial intervals. We found that in the bulk, likely due to the high laser power and the infinite nature of the periodic system, the IPA is inapplicable. These simulations similarly involve XC at the PBE level (similar level of theory to the cluster case), but where the geometries of the solutions were obtained as discussed above from MD simulations (with a different more accurate XC functional). At each supercell geometry (denoted snapshot, arising from the MD dynamical evolution of the liquid over picosecond timescales) the time-dependent current was extracted, (with labeling the angular orientation, and the snapshot geometry). We averaged the current evaluated from a total of six simulations at varying spatial orientations (different ) of the laser for a given snapshot (), and further for 50 snapshots (values ) separated by 5 ps each for a given solution concentration. From the averaged value of the current, , we computed the HHG spectrum by taking one time derivative and Fourier transforming with a temporal super gaussian filter : . This procedure converged the HHG yield at the first plateau region up to 19 eV and provided an isotropic odd-only harmonic response along the laser polarization axis, as expected from optical selection rules [80]. Further note that slightly lower laser intensities were employed in the bulk supercell to account for dielectric function changes that need not be taken into account in the finite cluster system. This reduced laser peak powers in actual simulations by a factor of 2.5. We employed simulations in supercells at a relative liquid density of 0.9 compared to standard density at atmospheric pressure and room temperature, because we noted that in typical conditions simulations did not reproduce the sharp harmonic suppression measured experimentally in Fig. 3. Based on this, we suggest that in the experimental geometry, the density of the liquid is likely lower than that at equilibrium due to the laser irradiation itself.
Quantum-orbit calculations with saddle-point approximation First, we solved the saddle-point equations in SFA with the approach introduced in [81]. With the saddle-point approximation, one can easily separate the long and short quantum paths, which have the opposite attochirp. In our calculations we only considered the short quantum paths due to their preferred phase matching. The calculated ionization and recollision times are both complex-valued. The real parts are shown in Figs. 2c and 4b and the imaginary parts are closely related to the ionization probability of the quantum orbit. For the simulations of the two-colour experiments, we considered the short quantum orbits from two optical cycles and summed their ionization probability amplitudes coherently. Their interference gives rise to the harmonic structures in the spectral domain.
Data Availability
The datasets generated and/or analysed during the current study will be published on the ETH data repository upon acceptance of the manuscript.
Code Availability
The Octopus package used for the TDDFT calculations is publicly available. The remaining computer codes are available from the corresponding authors on reasonable request.
Acknowledgements.
The authors thank Nirit Dudovich for insightful discussions, as well as Andreas Schneider and Mario Seiler for their contributions to the construction and improvements of the experiment. We acknowledge financial support from ETH Zürich. ON gratefully acknowledges support from the Young Faculty Award from the National Quantum Science and Technology program of Israel’s Council of Higher Education Planning and Budgeting Committee. G.T. gratefully acknowledges support from the Alexander von Humboldt Foundation. M.R. acknowledges funding by the European Union (ERC, QUADYMM, 101169761). Z.Y. acknowledges funding by JSPS KAKENHI Grant-in-Aid for Scientific Research (A) (No. 25H00864).Author Contributions
All authors participated in the discussion of the results and contributed to the manuscript.
Competing interests
The authors declare no competing interests.